### v1

# Eigenvalues of Heckman-Polychronakos operators

Abstract.
Heckman-Polychronakos operators form a prominent family of commuting differential–difference operators defined in terms of the Dunkl operators $\mathcal{D}_{i}$ as $\mathcal{P}_{m}=\sum_{i=1}^{N}(x_{i}\mathcal{D}_{i})^{m}$. They have been known since 1990s in connection with trigonometric Calogero-Moser-Sutherland Hamiltonian and Jack symmetric polynomials. We explicitly compute the eigenvalues of these operators for symmetric and skew-symmetric eigenfunctions, as well as partial sums of eigenvalues for general polynomial eigenfunctions.
1. Introduction
Our paper is motivated from two directions. From one side, the theory of quantum integrable systems or, more generally, spectral theory, seeks to construct families of commuting operators and understand the structure of their eigenvalues and eigenfunctions. From the other side, integrable probability often exploits operators to extract information about stochastic systems and their asymptotic behavior.
We deal with a family of differential–difference operators acting on polynomials in $N$ variables $x_{1},\dots,x_{N}$ and called the Heckman-Polychronakos operators, which were first introduced in [Hec91], [Pol92] in connection to the study of (trigonometric) Calogero-Moser-Sutherland Hamiltonian. The key properties of interest for these operators were that they commute with each other and with the CMS Hamiltonian, and that Jack symmetric polynomials (whose extensive theory was developed just a couple of years before that, see [Sta89], [Mac95, Chapter VI]) are their eigenfunctions.
Soon after these operators were introduced, two things happened. First, [Che91, Che94] introduced an alternative family of the Cherednik operators satisfying the same three properties, and, second, the theory was lifted up to the level of Macdonald polynomials (for whom Jack polynomials are a degeneration, cf. [Mac95, Chapter 6]). In subsequent years, Cherednik operators and their Macdonald versions were playing their roles in many developments, while Heckman-Polychronakos operators received much less attention, cf. [Cha24] for a recent historical overview of the theory. As a result, even the most basic properties of the latter operators, such as the formulas for eigenvalues, are still not well-understood by now. The goal of this paper is to fill in this gap in the literature and re-initiate the study of the Heckman-Polychronakos operators.
Some relationships between Cherednik and Heckman–Polychronakos operators are discussed in [SV15, Section 5]. Both families of operators can be constructed through Newton power sums of certain “basic” operators. However, these basic operators have different properties in two theories: in Cherednik theory the basic operators commute with each other, but their definition does not agree with the natural action of $S_{N}$ on $(x_{1},\dots,x_{N})$; in Heckman–Polychronakos theory the basic operators are defined in a $S_{N}$–equivariant way, but at the cost of not commuting with each other. [^1] We emphasize that although the basic operators do not commute, but the Heckman–Polychronakos operators do commute. The two parallel possible choices of basic operators in the Macdonald $(q,t)$–theory are further discussed in [NS19], see especially Section 2.4 there.
In the semiclassical limit [^2] This limit is related to degeneration of a Lie group into a Lie algebra and to an elementary transition $\lim\limits_{\varepsilon\to 0}(1+\varepsilon x)^{\varepsilon^{-1}y}=\exp(xy)$., Jack polynomials degenerate into multivariate Bessel functions (see e.g. [OO97, Section 4]) and the operators we study have a parallel degeneration to objects related to the rational (rather than trigonometric) Calogero-Moser-Sutherland Hamiltonian. Interestingly, Cherednik and Heckman–Polychronakos operators unite into a single family in this limit transition: the basic operators turn into the Dunkl operators of [Dun89] which both commute and are equivariant.
Switching to the integrable probability side, recently in [BGCG22, Xu23, GXZ24, KX24] the power sums of Dunkl operators were intensively used in the studies of $\beta$–ensembles of random matrix theory in problems related to additions of matrices and various limits of eigenvalue distributions. These random matrix distributions have natural discrete analogues [Ker00, BO05, GS15, DF16, BGG17, Mol23, DS19, GH19, Hua21, DK22, CDM23, DK24, GH24] which are intimately related to Jack polynomials. Many of the questions about matrix $\beta$–ensembles have versions for the discrete ensembles as well, and if one wants to attack them by the same techniques, then one should rely on a lifting of the power sums of Dunkl operators to the discrete setting, which would be either Cherednik or Heckman–Polychronakos operators. It is not yet clear, which of the two classes of operators would work better for the asymptotic problems, and one would like to be able to try both.
The desire to investigate the asymptotic questions about the stochastic systems from the previous paragraph was an initial motivation of our investigations of Heckman–Polychronakos operators in this paper, yet their theory turned out to be much more interesting than what we expected and, moreover, many open questions remain.
Acknowledgements. We would like to thank Evgeni Dimitrov and Grigori Olshanski for very helpful discussions. V.G. was partially supported by NSF grant DMS - 2246449.
2. Heckman–Polychronakos operators
Throughout the paper we fix $N=1,2,\dots$ and deal with the ring $\mathbb{C}_{N}=\mathbb{C}[x_{1},x_{2}\dots,x_{N}]$ of polynomials in $N$ variables. $\theta$ is an auxiliary parameter which enters into all our definitions: one can take it as a formal variable, or as a positive real number.
Definition 2.1.
For $1\leq i\leq N$ the Dunkl operator $\mathcal{D}_{i}$ is a linear operator in $\mathbb{C}_{N}$ acting by
(1) $\mathcal{D}_{i}=\frac{\partial}{\partial x_{i}}+\theta\sum_{j\neq i}\frac{1-(i,j)}{x_{i}-x_{j}},$
where $(i,j)$ stands for the operator permuting the variables $x_{i}$ and $x_{j}$.
Definition 2.2.
For $m\geq 1$ the Heckman-Polychronakos operator $\mathcal{P}_{m}$ is a linear operator in $\mathbb{C}_{N}$:
(2) $\mathcal{P}_{m}=\sum_{i=1}^{N}(x_{i}\mathcal{D}_{i})^{m},$
where $x_{i}$ stays for the operator of multiplication by $x_{i}$.
The operators $\mathcal{P}_{m}$ were first introduced and discussed in [Hec91], [Pol92]. The interest in them is based on three properties, whose proofs we recall in the Appendix:
  • $\mathcal{P}_{m}$ commute with each other as operators in $\mathbb{C}_{N}$: $\mathcal{P}_{m}\mathcal{P}_{k}=\mathcal{P}_{k}\mathcal{P}_{m}$.
  • When restricted onto the subring $\Lambda_{N}\subset\mathbb{C}_{N}$ of symmetric polynomials, the operator $\mathcal{P}_{2}$ is the (conjugated) Calogero-Moser-Sutherland Hamiltonian.
  • Symmetric Jack polynomials $J_{\lambda}(x_{1},\dots,x_{N};\,\theta)$, $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}\geq 0$, are eigenfunctions of all $\mathcal{P}_{m}$; see [Sta89], [Mac95, Chaprer VI] for general introduction to Jack polynomials and note that it uses the parameter $\alpha=1/\theta$.
In contrast to another famous family of operators satisfying these three properties — power sums of the (Jack versions of) Cherednik operators [Che91, Che94] – the individual terms $(x_{i}\mathcal{D}_{i})^{m}$ do not commute. [^3] For a fruitful study of commutators of $\mathcal{D}_{i}$ and $x_{j}$ operators see [FH15]. On the other hand, the advantage of Heckman-Polychronakos operators is that their definition is simpler than the Cherednik ones and they are equivariant with respect to the symmetric group action.
The aim of this paper is to study the spectrum of the operators $\mathcal{P}_{m}$: we would like to understand the eigenvalues corresponding to Jack polynomials $J_{\lambda}$, as well as those corresponding to other non-symmetric eigenfunctions.
The $m=1$ operator is simpler than others:
$\sum_{i=1}^{N}x_{i}\mathcal{D}_{i}=\sum_{i=1}^{N}x_{i}\frac{\partial}{\partial x_{i}}+\theta\sum_{1\leq i<j\leq N}\frac{x_{i}-x_{j}}{x_{i}-x_{j}}[1-(i,j)]=\sum_{i=1}^{N}x_{i}\frac{\partial}{\partial x_{i}}+\frac{\theta N(N-1)}{2}-\theta\sum_{1\leq i<j\leq N}(i,j).$
The appearance of $\sum_{i<j}(i,j)$ operator hints on the relevance of the representation theory of $S_{N}$ for the computations of eigenvalues and we will see that this is indeed the case.
3. Symmetric eigenfunctions
In this section we compute the eigenvalues corresponding to the (symmetric) Jack polynomials.
A partition of $n$ is a sequence of non-negative integers $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$, such that $\sum_{i}\lambda_{i}=n$. One can write $\lambda\vdash n$ or $|\lambda|=n$ in this situation. The length of $\lambda$, denoted $\ell(\lambda)$ is the number of non-zero parts in $\lambda$. In particular, if $\ell(\lambda)\leq N$, then $\lambda_{N+1}=0$ and we can represent $\lambda$ as $N$–tuple $(\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}\geq 0)$.
We also recall the notation $h_{m}(x_{1},\dots,x_{N})$ for complete homogeneous symmetric functions and introduce their relatives, see [Mac95, Example 19, Section I.2] for similar (but slightly different) functions:
(3) $h^{(r)}_{m}(x_{1},\dots,x_{N})=\sum_{r-\text{element subsets }\{i_{1},\dots,i_{r}\}\subset\{1,2,\dots,N\}}h_{m}(x_{i_{1}},\dots,x_{i_{r}}).$
For $m<0$ we set $h^{(r)}_{m}=0$. Clearly,
$h^{(1)}_{m}(x_{1},\dots x_{N})=x_{1}^{m}+x_{2}^{m}+\dots+x_{N}^{m},$
$h^{(2)}_{m-1}(x_{1},\dots,x_{N})=\sum_{i<j}\frac{x_{i}^{m}-x_{j}^{m}}{x_{i}-x_{j}},$
and
$h^{(N)}_{m}(x_{1},\dots,x_{N})=h_{m}(x_{1},\dots,x_{N})=\sum_{j_{1}+j_{2}+\dots+j_{N}=m}x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{N}^{j_{N}}.$
Theorem 3.1.
Denote $\ell_{i}=\lambda_{i}+\theta(N-i)$, $1\leq i\leq N$. For each $\lambda_{1}\geq\dots\geq\lambda_{N}\geq 0$ we have
(4) $\mathcal{P}_{m}J_{\lambda}(x_{1},\dots,x_{N};\,\theta)=\mathrm{eig}_{m}(\lambda)J_{\lambda}(x_{1},\dots_{N};\,\theta),$
where
(5) $\mathrm{eig}_{m}(\lambda)=h^{(1)}_{m}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}-\theta h^{(2)}_{m-1}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}+\dots+(-\theta)^{N-1}h^{(N)}_{m+1-N}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}.$
Theorem 3.1 is a combination of Theorem 5.4 and Lemma 5.5 below. An equivalent compact expression for the eigenvalue is:
(6) $\mathrm{eig}_{m}(\lambda)=\begin{pmatrix}1&1&1&\ldots&1\end{pmatrix}\begin{pmatrix}\ell_{1}&-\theta&-\theta&\ldots&-\theta\\ 0&\ell_{2}&-\theta&\ldots&-\theta\\ 0&0&\ell_{3}&\ldots&-\theta\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&\ell_{N}\end{pmatrix}^{m}\begin{pmatrix}1\\ 1\\ 1\\ \vdots\\ 1\end{pmatrix}.$
The eigenvalues $\mathrm{eig}_{m}(\lambda)$ also have a simple generating function:
Proposition 3.2.
Denoting $\ell_{i}=\lambda_{i}+\theta(N-i)$, for each $\lambda_{1}\geq\dots\geq\lambda_{N}\geq 0$ we have:
(7) $1-\theta z\sum_{m=0}^{\infty}\mathrm{eig}_{m}(\lambda)z^{m}=\prod_{i=1}^{N}\left(1-\frac{\theta z}{1-\ell_{i}z}\right)=\prod_{i=1}^{N}\frac{1-(\ell_{i}+\theta)z}{1-\ell_{i}z}.$
Proof.
For each $1\leq r\leq N$, using the generating function for the complete homogeneous functions $h_{m}$ (see [Mac95, Chapter 1, Section 2]) we have
$\sum_{m=0}^{\infty}h^{(r)}_{m+1-r}(\ell_{1},\dots,\ell_{N})z^{m}=z^{r-1}\sum_{r-\text{element subsets }\{i_{1},\dots,i_{r}\}\subset\{1,2,\dots,N\}}\sum_{k=0}^{\infty}h_{k}(\ell_{i_{1}},\dots,\ell_{i_{r}})z^{k}\\ =z^{r-1}\sum_{r-\text{element subsets }\{i_{1},\dots,i_{r}\}\subset\{1,2,\dots,N\}}\prod_{j=1}^{r}\frac{1}{1-\ell_{i_{j}}z}.$
Hence, plugging into (5) we get
$1-\theta z\sum_{m=0}^{\infty}\mathrm{eig}_{m}(\lambda)z^{m}=\sum_{\begin{smallmatrix}\mathcal{A}\subset\{1,\dots,N\}\end{smallmatrix}}(-\theta z)^{|\mathcal{A}|}\prod_{a\in\mathcal{A}}\frac{1}{1-\ell_{a}z}=\prod_{i=1}^{N}\left(1-\frac{\theta z}{1-\ell_{i}z}\right).\qed$
Remark 3.3.
The fact that symmetric Jack polynomials are eigenfunctions of $\mathcal{P}_{m}$ is not new: there is no doubt that the authors of [Hec91], [Pol92], or [LV96] were aware of it. However, the formulas for the eigenvalues are more elusive and we could not locate (5), (6), or (7) in the literature.
G. Olshanski pointed to us that in the special case [^4] At $\theta=1$ Jack polynomials turn into Schur functions, which are characters of the irreducible representations of the unitary group $U(N)$. $\theta=1$, (6) matches the formula for the eigenvalue of the classical Casimir operator $C_{m}$, as given in [Zhe73, (*) and Theorem 2 in Section 60].
Formula (7) is similar to the expression for eigenvalues of Nazarov-Sklyanin operators in the space of polynomials in infinitely many variables, see [NS13, Section 6].
4. Classification of eigenfunctions
Outside Jack polynomials, other eigenfunctions of $\mathcal{P}_{m}$ are much less understood, and in this section we propose a classification for them.
We need the dominance order on partitions:
Definition 4.1.
Let $\lambda=(\lambda_{1}\geq\dots\lambda_{N}\geq 0)$ and $\tilde{\lambda}=(\tilde{\lambda}_{1}\geq\dots\geq\tilde{\lambda}_{N}\geq 0)$ be two distinct partitions of length at most $N$. We write $\lambda\succ\tilde{\lambda}$, if $\sum_{i=1}^{N}\lambda_{i}=\sum_{i=1}^{N}\tilde{\lambda}_{i}$ and for all $1\leq k\leq N$, we have $\sum_{i=1}^{k}\lambda_{i}\geq\sum_{i=1}^{k}\tilde{\lambda}_{i}$.
Let $\gamma=(\gamma_{1},\dots,\gamma_{N})$ be a degree sequence of a monomial in $x_{1},\dots,x_{N}$ and write $x^{\gamma}$ for $x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}}\cdots x_{N}^{\gamma_{N}}$. We let $\gamma^{+}$ denote the partition $\lambda_{1}\geq\dots\lambda_{N}\geq 0$ obtained by rearranging $\gamma_{i}$ in nonincreasing order.
Definition 4.2.
For $\lambda=(\lambda_{1}\geq\dots\lambda_{N}\geq 0)$, we let $V_{\lambda}$ denote the linear space spanned by all monomials $x^{\gamma}$ with $\gamma^{+}=\lambda$.
Note that $V_{\lambda}$ has a structure of a representation of symmetric group $S_{N}$ permuting the variables $x_{i}$. We write $(i,j)$ for the transpositions in the symmetric group and $(i,j,k\dots)$ for one-cycle permutations. When we write $\sigma_{2}\sigma_{1}f$ for a polynomial $f(x_{1},\dots,x_{N})$ and permutations $\sigma_{1},\sigma_{2}\in S_{N}$, we mean that we first permute the variables according to $\sigma_{1}$ and then permute according to $\sigma_{2}$. For instance:
$(1,2)\,x_{1}^{2}x_{2}=x_{1}x_{2}^{2},\qquad(1,3)(1,2)\,x_{1}^{2}x_{2}=(1,2,3)\,x_{1}^{2}x_{2}=x_{2}^{2}x_{3}.$
Let $V_{\lambda;\tau}$ denote the isotypical component of the irreducible representation $\tau$ in $V_{\lambda}$, so that
(8) $V_{\lambda}=\bigoplus_{\tau}V_{\lambda;\tau}.$
We recall that irreducible representations of symmetric group $S_{N}$ are parameterized by partitions $\tau\vdash N$, see [Mac95, Section I.7], [OV96], [Jam06] for the reviews of the representation theory of $S_{N}$. If we denote $\mathrm{Mult}(\lambda)$ the partition of $N$ representing multiplicities of coordinates in $\lambda$ (for each $a\geq 0$ we count how many times $a$ appears in $\lambda$, and then reorder decreasingly the resulting numbers), then the summation in (8) goes over $\tau$ such that $\mathrm{Mult}(\lambda)\preceq\tau$, see, e.g. [Mac95, Remark on page 115].
Outside several special cases, for generic $\lambda$ and $\tau$ the numbers $\dim V_{\lambda;\tau}$ do not admit a closed formula, although they are of interest in algebraic combinatorics (being equivalent to Kostka numbers). It is straightforward to show that the dependence of $\dim V_{\lambda;\tau}$ on $\lambda$ is only through $\mathrm{Mult}(\lambda)$.
Theorem 4.3.
For each triplet $(\lambda,\tau,i)$ with $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{N}\geq 0)$, $\tau\vdash N$, and $1\leq i\leq\dim V_{\lambda;\tau}$, the operators $\mathcal{P}_{m}$, $m\geq 1$, have a joint eigenfunction $F_{\lambda,\tau,i}$, such that:
  1. (1)
    $F_{\lambda,\tau,i}$ is a homogeneous polynomial of degree $|\lambda|$.
  2. (2)
    All monomials $x^{\gamma}$ with non-zero coefficients in $F_{\lambda,\tau,i}$ are such that $\gamma^{+}\preceq\lambda$.
  3. (3)
    The part of $F_{\lambda,\tau,i}$ spanned by monomials in $V_{\lambda}$ belongs to $V_{\lambda;\tau}$.
The functions $F_{\lambda,\tau,i}$ span $\mathbb{C}_{N}$ and there are no other polynomial eigenfunctions.
Remark 4.4.
We do not propose any explicit form for the dependence of $F_{\lambda,\tau,i}$ on $i$. One reason is that the combination of operators $\mathcal{P}_{m}$ and the structure of $S_{N}$–representation is not sufficient to distinguish them. Indeed, for any coefficients $\alpha$ and $\beta$ and any $g\in S_{N}$, the function $\alpha\cdot F_{\lambda,\tau,i}+\beta\cdot g\circ F_{\lambda,\tau,i}$ is an eigenfunction of $\mathcal{P}_{m}$ satisfying the same three properties and with the same eigenvalue.
Examples 4.5.
For each $\lambda$, one can choose $\tau$ to be a one-part partition, $\tau=(N)$, corresponding to trivial representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the sum of all monomials in $V_{\lambda}$. The corresponding eigenfunction is the symmetric Jack polynomial.
For each $\lambda$ with distinct parts, one can choose $\tau$ to be a unique $N$–parts partition $\tau=1^{N}$, corresponding to the sign representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the signed sum of all monomials in $V_{\lambda}$, i.e. $\sum_{\sigma\in S_{N}}(-1)^{\sigma}\circ x_{\sigma(1)}^{\lambda_{1}}x_{\sigma(2)}^{\lambda_{2}}\cdots x_{\sigma(N)}^{\lambda_{N}}$. The corresponding eigenfunction is a skew-symmetric polynomial, which can be expressed as alternating sum of all non-symmetric Jack polynomials labeled by $N!$ compositions obtained by permuting the coordinates of $\lambda$ or, alternatively, as the symmetric Jack polynomial with shifted $\theta$ and multiplied by the $\prod_{i<j}(x_{i}-x_{j})$, see [BF97, (2.40)].
For $\lambda=(a,a,\dots,a)$, the space $V_{\lambda}$ is one-dimensional and only $\tau=(N)$ is possible.
For $\lambda$ with distinct parts, the space $V_{\lambda}$ is isomorphic to regular representation of $S_{N}$ of dimension $N!$. For each partition $\tau\vdash N$, $\dim V_{\lambda;\tau}$ is the square of the dimension of irreducible representation $\tau$, and this is the number of eigenfunctions $F_{\lambda,\tau,i}$.
In the rest of this section we prove Theorem 4.3. Two ingredients of the proof are triangularity of Lemma 4.6 and self-adjointness of Lemma 4.7.
Let us introduce a linear operator $T_{i}$, which acts in $\mathbb{C}_{N}$ and leaves each $V_{\lambda}$ invariant:
(9) $T_{i}x^{\gamma}=\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)x^{\gamma}-\theta\sum_{j\mid\gamma_{j}>\gamma_{i}}(i,j)x^{\gamma},\qquad 1\leq i\leq N.$
Lemma 4.6.
For any partition $\lambda$ and any degree sequence $\gamma$ with $\gamma^{+}=\lambda$, we have
$\mathcal{P}_{m}[x^{\gamma}]=(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})[x^{\gamma}]+(\text{linear combination of monomials }x^{\tilde{\gamma}}\text{ with }\tilde{\gamma}^{+}\prec\lambda).$
Proof.
Note that
(10) $x_{i}\frac{1-(i,j)}{x_{i}-x_{j}}\bigl{[}x_{1}^{\gamma_{1}},\dots,x_{N}^{\gamma_{N}}\bigr{]}=\prod_{a\neq i,j}x_{a}^{\gamma_{a}}\cdot x_{i}\frac{x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}-x_{i}^{\gamma_{j}}x_{j}^{\gamma_{i}}}{x_{i}-x_{j}}\\ =\prod_{a\neq i,j}x_{a}^{\gamma_{a}}\cdot\begin{cases}x_{i}^{\gamma_{i}}x_{j}^{\gamma_{j}}+x_{i}^{\gamma_{i}-1}x_{j}^{\gamma_{j}+1}+\dots+x_{i}^{\gamma_{j}+1}x_{j}^{\gamma_{i}-1},&\gamma_{j}<\gamma_{i},\\ -x_{i}^{\gamma_{j}}x_{j}^{\gamma_{i}}-x_{i}^{\gamma_{j}-1}x_{j}^{\gamma_{i}+1}-\dots-x_{i}^{\gamma_{i}+1}x_{j}^{\gamma_{j}-1},&\gamma_{j}>\gamma_{i},\\ 0,&\gamma_{j}=\gamma_{i}.\end{cases}$
Hence,
$x_{i}\mathcal{D}_{i}[x^{\gamma}]=T_{i}x^{\gamma}+(\text{linear combination of monomials }x^{\tilde{\gamma}}\text{ with }\tilde{\gamma}^{+}\prec\gamma^{+}).$
Iterating the last identity for $\mathcal{P}_{m}=\sum_{i=1}^{N}(x_{i}\mathcal{D}_{i})^{m}$, we arrive at the statement of the lemma. ∎
Lemma 4.7.
Consider a scalar product on $\mathbb{C}_{N}$, where $\overline{g}$ stays for complex conjugation:
(11) $\langle f,g\rangle=\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}f\bigl{(}e^{\mathbf{i}\phi_{1}},\dots,e^{\mathbf{i}\phi_{N}}\bigr{)}\overline{g\bigl{(}e^{\mathbf{i}\phi_{1}},\dots,e^{\mathbf{i}\phi_{N}}\bigr{)}}\prod_{1\leq i<j\leq N}\bigl{|}e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}\bigr{|}^{2\theta}\,\mathrm{d}\phi_{1}\dots\mathrm{d}\phi_{N}.$
Then the operators $x_{i}\mathcal{D}_{i}$, $1\leq i\leq N$, are self-adjoint with respect to it: for any $f,g,\in\mathbb{C}_{N}$
$\langle x_{i}\mathcal{D}_{i}f,g\rangle=\langle f,x_{i}\mathcal{D}_{i}g\rangle.$
Proof.
Versions of this statement can be found in [Dun91] and [Hec91]. Changing the variables $x_{i}=\exp(-\mathbf{i}\phi_{i})$, the operator $x_{i}\mathcal{D}_{i}$ is transformed into
$\mathbf{i}\frac{\partial}{\partial\phi_{i}}+\theta\sum_{j\neq i}\frac{1-(i,j)}{1-e^{\mathbf{i}(\phi_{j}-\phi_{i})}},$
and we need to check that it is self-adjoint with respect to the scalar product
(12) $\langle f,g\rangle_{\phi}=\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}f\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}\overline{g\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}}\prod_{1\leq i<j\leq N}\bigl{|}e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}\bigr{|}^{2\theta}\,\mathrm{d}\phi_{1}\dots\mathrm{d}\phi_{N}.$
We first check that the operator $\frac{(i,j)}{1-\exp(\mathbf{i}(\phi_{j}-\phi_{i}))}$ is self-adjoint, which is the identity
$\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}\frac{\bigl{[}(i,j)f\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}\bigr{]}\overline{g\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}}}{1-e^{\mathbf{i}(\phi_{j}-\phi_{i})}}\prod_{1\leq i<j\leq N}\bigl{|}e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}\bigr{|}^{2\theta}\,\mathrm{d}\phi_{1}\dots\mathrm{d}\phi_{N}\\ \stackrel{{\scriptstyle?}}{{=}}\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}\frac{f\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}\bigl{[}(i,j)\overline{g\bigl{(}\phi_{1},\dots,\phi_{N}\bigr{)}}\bigr{]}}{1-e^{-\mathbf{i}(\phi_{j}-\phi_{i})}}\prod_{1\leq i<j\leq N}\bigl{|}e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}\bigr{|}^{2\theta}\,\mathrm{d}\phi_{1}\dots\mathrm{d}\phi_{N},$
that is tautologically true by renaming $\phi_{i}\leftrightarrow\phi_{j}$ in the integral. It remains to check that $\mathbf{i}\frac{\partial}{\partial\phi_{i}}+\theta\sum_{j\neq i}\frac{1}{1-\exp(\mathbf{i}(\phi_{j}-\phi_{i}))}$ is self-adjoint. For that we record the fact that $\frac{\partial}{\partial\phi_{i}}$ derivative of the integrand in the right-hand side of (12) integrates to zero:
$0=\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}\left[\frac{\partial}{\partial\phi_{i}}f\overline{g}+f\frac{\partial}{\partial\phi_{i}}\overline{g}+\theta f\overline{g}\sum_{j\neq i}\frac{\mathbf{i}e^{\mathbf{i}\phi_{i}}}{e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}}+\theta f\overline{g}\sum_{j\neq i}\frac{-\mathbf{i}e^{-\mathbf{i}\phi_{i}}}{e^{-\mathbf{i}\phi_{i}}-e^{-\mathbf{i}\phi_{j}}}\right]\\ \times\prod_{1\leq i<j\leq N}\bigl{(}e^{\mathbf{i}\phi_{i}}-e^{\mathbf{i}\phi_{j}}\bigr{)}^{\theta}\bigl{(}e^{-\mathbf{i}\phi_{i}}-e^{-\mathbf{i}\phi_{j}}\bigr{)}^{\theta}\,\mathrm{d}\phi_{1}\dots\mathrm{d}\phi_{N}.$
Multiplying by $\mathbf{i}$ we get the desired self-adjointness statement. ∎
Remark 4.8.
Following [Dun91, Section 3] and [Dun08, Section 3], there is another scalar product, which also makes $x_{i}\mathcal{D}_{i}$ self-adjoint. This other scalar product is defined through $\langle f,g\rangle=f(\mathcal{D}_{1},\dots,\mathcal{D}_{N})g(x_{1},\dots,x_{N})\bigr{|}_{x_{1}=\dots=x_{N}=0}$ and it makes $x_{i}$ the adjoint operator to $\mathcal{D}_{i}$.
Proof of Theorem 4.3.
Since operators $\mathcal{P}_{m}$ are self-adjoint by Lemma 4.7, commutative by Lemma 7.1 in the appendix, and essentially finite-dimensional (because $\mathcal{P}_{m}$ preserves the space of degree $k$ polynomials for each $k$), they have a joint eigenbasis and we only need to show that the eigenfunctions satisfy the claimed properties. Consider the space $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. By Lemma 4.6, this is an invariant space for $\mathcal{P}_{m}$. Because it is finite-dimensional and $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, this space has a complete system of eigenfunctions of $\mathcal{P}_{m}$. Inductively counting eigenfunctions, which belong to such spaces with smaller $\lambda$, we conclude that in $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ there are precisely $\dim V_{\lambda}$ eigenfunctions, which have at least one non-zero coefficient among monomials from $V_{\lambda}$: these are various $F_{\lambda,\tau,i}$ with fixed $\lambda$ and varying $\tau$ and $i$. By triangularity of Lemma 4.6, for each such function, its part spanned by the monomials from $V_{\lambda}$ is necessary an eigenfunction of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. The latter operator has a symmetric expression in terms of $x_{1},\dots,x_{N}$, and therefore it commutes with the action of $S_{N}$. Hence, by the Schur’s lemma, each $V_{\lambda;\tau}$ is an invariant subspace for $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. Therefore, there are precisely $\dim V_{\lambda;\tau}$ eigenfunctions of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$ inside $V_{\lambda;\tau}$. By triangularity, each of them has a unique extension to an eigenfunction of $\mathcal{P}_{m}$ and these are $F_{\lambda,\tau,i}$, $1\leq i\leq\dim V_{\lambda;\tau}$. ∎
5. Sums of eigenvalues
In general, the eigenvalue of the operator $\mathcal{P}_{m}$ on the function $F_{\lambda,\tau,i}$ can be expressed as a root of a polynomial equation (of degree at most $\dim(\tau)$) with coefficients being polynomials in $\lambda_{i}$ and $\theta$; therefore, this eigenvalue can be quite complicated, see, e.g. (33) in Section 6. We found that the sums of all eigenvalues corresponding to the same $\lambda$ and $\tau$, are much more manageable.
5.1. Distinct degrees of variables
For an irreducible representation of $S_{N}$ labeled by $\tau\vdash N$, we let $\dim\tau$ be its dimension and let $\chi^{\tau}(g)$ be its character. In particular, $\dim(\tau)=\chi^{\tau}(\mathrm{Id})$. We also recall the notation $h_{m}^{(r)}$ from (3).
Theorem 5.1.
Let $\mathrm{eig}_{m}(\lambda,\tau,i)$ denote the eigenvalue of $\mathcal{P}_{m}$ on the eigenfunction $F_{\lambda,\tau,i}$ of Theorem 4.3. Suppose that $\lambda$ has all distinct parts. Then, setting $\ell_{i}=\lambda_{i}+\theta(N-i)$, we have
(13) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=\dim\tau\sum_{k=1}^{\min(m+1,N)}(-\theta)^{k-1}h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}\chi^{\tau}\bigl{(}(1,2,\dots,k)),$
where $(1,2,\dots,k)$ is the permutation having a single $k$-cycle and $N-k$ fixed points.
When $\tau=(N)$, $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into (5); this eigenvalue corresponds to the symmetric eigenfunction. When $\tau=1^{N}$, also $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into a formula similar to (5) but with $(-\theta)^{k-1}$ replaced with $\theta^{k-1}$; this eigenvalue corresponds to a skew-symmetric eigenfunction.
In general, the numbers $\chi^{\tau}\bigl{(}(1,2,\dots,k))$ admit a somewhat explicit formula in terms of $\tau=(\tau_{1}\geq\dots\geq\tau_{N}\geq 0)$, see, e.g., [Mac95, Example I.7.7] for the following expression and [IO02, Section 3] for some others
(14) $\chi^{\tau}\bigl{(}(1,2,\dots,k))=\dim\tau\frac{(N-k)!}{N!}\sum_{i=1}^{N}\frac{(\tau_{i}+N-i)!}{(\tau_{i}+N-i-k)!}\prod_{j\neq i}\frac{\tau_{i}-i-\tau_{j}+j-k}{\tau_{i}-i-\tau_{j}+j}.$
If parts of $\lambda$ are allowed to coincide, (13) is replaced by a significantly more complicated formula, which we present in Theorem 5.4 below.
For the proof of Theorem 5.1 we use $T_{i}$ of (9). We also need an additional object from representation theory of $S_{N}$. For each irreducible representation $\tau$ of $S_{N}$, denote
(15) $\pi_{\tau}=\frac{\dim\tau}{N!}\sum_{g\in S_{N}}\chi^{\tau}(g)g.$
[Ser77, Theorem 8] says that in each representation of $S_{N}$, $\pi_{\tau}$ acts as an orthogonal projector (with respect to an $S_{N}$–invariant scalar product) onto isotypical component of $\tau$.
Lemma 5.2.
The left-hand side of (13) can be evaluated as
(16) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}(T_{1}^{m}+\dots+T_{N}^{m})\right]=N\,\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right].$
Proof.
The second equality in (16) follows from the invariance of both the Trace and $\pi_{\tau}$ under conjugations with $g\in S_{N}$, and we only prove the first one.
Recall from the proof of Theorem 4.3 that $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of $(T_{1}^{m}+\dots+T_{n}^{m})$, corresponding to eigenfunctions in $V_{\lambda;\tau}$. Hence, $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of the restriction of $(T_{1}^{m}+\dots+T_{n}^{m})$ onto $V_{\lambda;\tau}$, which is the same as eigenvalues of $\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}$. The sum of the eigenvalues coincides with trace, and we get $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}\right]$. Since $\mathrm{Trace}[AB]=\mathrm{Trace}[BA]$, we can move one $\pi_{\tau}$ inside the trace from right to the left, and then using the projecting property $\pi_{\tau}\pi_{\tau}=\pi_{\tau}$, we arrive at (16). ∎
Proof of Theorem 5.1.
We use Lemma 5.2 and compute $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$ combining (9) with (15). The computation is based on the observation that $V_{\lambda}$ is isomorphic to the regular representation of $S_{N}$, and therefore in the trace computation all products of $g$ and $(1,j)$ not equal to the identity permutation do not give any contribution.
Let us compute $(T_{1})^{m}x^{\gamma}$ for $\gamma=(\gamma_{1},\gamma_{2},\dots,\gamma_{N})$. We have
(17) $(T_{1})^{m}x^{\gamma}=\sum_{k=1}^{m+1}\sum_{j_{1},\dots,j_{k-1}}p_{k;q}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]\cdot(-\theta)^{k-1}(1,j_{k-1})(1,j_{k-2})\cdots(1,j_{1})x^{\gamma},$
where $j_{1},\dots,j_{k-1}$ are distinct indices, such that each of them corresponds to distinct $\gamma_{j_{k-1}}>\gamma_{j_{k-2}}>\dots>\gamma_{j_{1}}>\gamma_{1}$. The transpositions $(1,j_{a})$ correspond to the second term in (9) and the ordering appears because the operator $T_{1}$ can only increase the degree of $x_{1}$, but never decrease. On the other hand, the prefactor $p_{k;q}$ collects all the factors coming from the first term; it is a polynomial in various $\theta$-shifted $\gamma_{i}$ of degree $(m+1-k)$, which depends on the collection $\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}$ and on $\theta$. Note also that the product of $k-1$ transpositions becomes a $k$–cycle
$(1,j_{k-1})(1,j_{k-2})\cdots(1,j_{1})=(1,j_{1},j_{2},\dots,j_{k-1}).$
We further represent $g$ in (15) as
$g=\tilde{g}\cdot(1,j_{1})\cdots(1,j_{k-2})(1,j_{k-1}),$
and note that when we plug into $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$, and multiply the sums, only the terms with $\tilde{g}=\mathrm{Id}$ give non-zero contributions. We also note that the trace is a central function and its values on all $k$–cycles are the same. Therefore,
(18) $N\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]=\frac{\dim\tau}{(N-1)!}\sum_{k=1}^{m+1}(-\theta)^{k-1}\chi^{\tau}\bigl{(}(1,2,\dots,k)\bigr{)}\sum_{\gamma}\sum_{j_{1},\dots,j_{k-1}}p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}].$
It remains to compute the last double sum and we claim that
(19) $\frac{1}{(N-1)!}\sum_{\gamma}\sum_{j_{1},\dots,j_{k-1}}p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]\\ =\text{Coefficient of }(-\theta)^{k-1}\text{ in: }\begin{pmatrix}1&1&\dots&1\end{pmatrix}\cdot\begin{pmatrix}\ell_{1}&-\theta&\dots&&&-\theta\\ 0&\ell_{2}&-\theta&&&-\theta\\ 0&0&\ell_{3}&-\theta&&\vdots\\ \vdots&\vdots&\vdots&&&-\theta\\ 0&0&&\dots&0&\ell_{N}\end{pmatrix}^{m}\cdot\begin{pmatrix}1\\ 1\\ \vdots\\ 1\end{pmatrix}$
Indeed, there are $N!$ terms in the sum over $\gamma$, however, only the rank of $\gamma_{1}$ among all coordinates of $\gamma$ matters for the computation of the following $\sum_{j_{1},\dots,j_{k-1}}$. Hence, $\frac{1}{(N-1)!}\sum_{\gamma}$ can be interpreted as a sum with $N$ terms corresponding to $N$ possible ranks of $\gamma_{1}$.
Next, we notice that the first term in (9) is always equal to of $\ell_{a}$ and therefore matches the diagonal element of the matrix in (19). Similarly, the $-\theta(i,j)$ terms in (9) match the off-diagonal $(-\theta)$ elements in the matrix in (19). Hence, we see that the combinatorics of rising the matrix to the $m$–th power in (19) and raising $T_{i}$ to the $m$–th power is exactly the same. Hence, the identity (19).
It remains to identify the right-hand side of (19) with $h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}$, which is immediate from the definition (3): the $k$–element subset $\{i_{1},\dots,i_{k}\}$ encodes those diagonal elements of the matrix in (19), which enter into the computation of the $m$–th power.
Finally, the summation $\sum_{k=1}^{m+1}$ in (18) can be restricted to $\sum_{k=1}^{\min(m+1,N)}$ in (13), because for $k>N$ finding $k$ distinct coordinates $\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}$ becomes impossible and the corresponding sum in (18) is empty. ∎
5.2. General degrees
We now proceed to the most general result, in which parts of $\lambda$ are allowed to have multiplicities.
Given an integer $0\leq p\leq N$, a multiplicity composition (i.e. a sequence of $p$ positive integers) $\mathbf{n}=(n_{1},n_{2},\dots,n_{p}>0)$ with $\sum_{i=1}^{p}n_{i}=N$, and a degree sequence $\mathbf{d}=(d_{1}>d_{2}>\dots>d_{p}\geq 0)$, we introduce a partition
(20) $\lambda=\underbrace{d_{1},\dots,d_{1}}_{n_{1}},\underbrace{d_{2},\dots,d_{2}}_{n_{2}},\dots,\underbrace{d_{p},\dots,d_{p}}_{n_{p}}.$
For any $k$–element subset $\mathcal{A}\subset\{1,\dots,p\}$ of the form $\{a_{1}<a_{2}<\dots<a_{k}\}$, we introduce averaged characters also called “spherical functions”:
(21) $\chi^{\tau}[\mathcal{A};\mathbf{n}]=\frac{1}{(n_{1})!(n_{2})!\cdots(n_{p})!}\sum_{\tilde{g}\in S_{n_{1}}\times S_{n_{2}}\times\dots\times S_{n_{p}}}\,\,\chi^{\tau}(\tilde{g}\cdot\mathfrak{c}),$
where $\mathfrak{c}$ is a $k$-cycle, joining together the $S_{n_{a_{i}}}$, $1\leq i\leq k$, subgroups, in the order of increasing $i$. More explicitly:
$\mathfrak{c}=(n_{1}+\dots+n_{a_{1}}-1,\,n_{1}+\dots+n_{a_{2}}-1,\,\dots,n_{1}+\dots+n_{a_{k-1}}-1,\,n_{1}+\dots+n_{a_{k}}-1).$
Remark 5.3.
Due to centrality of the characters, one can also choose other numbers representing $S_{n_{a_{i}}}$ subgroups when defining $\mathfrak{c}$ and arrive at exactly the same function $\chi^{\tau}[\mathcal{A};\mathbf{n}]$. For instance, $n_{1}+\dots+n_{a_{1}}-1$ can be replaced by any $j$, such that $n_{1}+\dots+n_{a_{1}-1}<j\leq n_{1}+\dots+n_{a_{1}}-1$.
We also introduce the shifted degrees:
~$\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p}),\qquad 1\leq i\leq p.$
Note that {~$\{\tilde{\ell}_{i}\}_{i=1}^{p}$ is a subset of $\{\ell_{i}=\lambda_{i}+\theta(N-i)\}_{i=1}^{N}$. Finally, we define complete homogeneous functions in subsets of these numbers:
(22) $h^{\mathcal{A}}_{m}=h_{m}(\tilde{\ell}_{a_{1}},\tilde{\ell}_{a_{2}},\dots,\tilde{\ell}_{a_{k}}),\qquad\mathcal{A}=\{a_{1},a_{2},\dots,a_{k}\}.$
Note that for $p=N$, the function $h^{(k)}_{m}$, as in (3), is the sum of $h^{\mathcal{A}}$ over all $k$–element subsets.
Theorem 5.4.
Let $\mathrm{eig}_{m}(\lambda,\tau,i)$ denote the eigenvalue of $\mathcal{P}_{m}$ on the eigenfunction $F_{\lambda,\tau,i}$ of Theorem 4.3. Suppose that $\lambda$ has the general form (20). Then, using the notations (21) and (22), we have
(23) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=\dim\tau\sum_{k=1}^{\min(m+1,p)}(-\theta)^{k-1}\sum_{\begin{smallmatrix}\mathcal{A}\subset\{1,2\dots,p\}\\ |\mathcal{A}|=k\end{smallmatrix}}h^{\mathcal{A}}_{m+1-k}\cdot\chi^{\tau}[\mathcal{A};\mathbf{n}]\cdot\prod_{a\in\mathcal{A}}n_{a}.$
If $\lambda$ has distinct parts, i.e. $n_{1}=n_{2}=\dots=n_{N}=1$, then there is no averaging in (21) and $\chi^{\tau}[\mathcal{A};\mathbf{n}]=\chi^{\tau}\bigl{(}(12\dots k)\bigr{)}$, where $k=|\mathcal{A}|$. Hence, Theorem 5.4 matches Theorem 5.1.
If $\tau=(N)$ is the trivial representation, then the eigenfunction is the Jack polynomials (see Lemma 7.4 for some details) and simultaneously $\chi^{\tau}[\mathcal{A};\mathbf{n}]=1$. Yet, the expression we get in (23) is slightly different from the one of Theorem 3.1, and we prove their equivalence in Lemma 5.5 below.
If $\tau=(1^{N})$ is the sign representation, then $\chi^{\tau}[\mathcal{A};\mathbf{n}]=0$, unless $\lambda$ has all distinct parts. In fact, there is no $(1^{N})$–type component in $V_{\lambda}$ for $\lambda$ with non-trivial multiplicities of parts.
Explicit evaluation of $\chi^{\tau}[\mathcal{A};\mathbf{n}]$ for $\tau=(N-1,1)$ is later given in Theorem 5.6 and the $p=2$ case is explained in Corollary 5.7.
Proof of Theorem 5.4.
The proof is an extended version of the argument in Theorem 5.1. Let us compute $(T_{1})^{m}x^{\gamma}$ for $\gamma^{+}=\lambda$, just as in (17), but now allowing the coordinates of $\gamma_{i}$ to coincide. The answer has the same form:
(24) $(T_{1})^{m}x^{\gamma}=\sum_{k=1}^{m+1}\sum_{j_{1},\dots,j_{k-1}}p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]\cdot(-\theta)^{k-1}(1,j_{k-1})(1,j_{k-2})\cdots(1,j_{1})x^{\gamma},$
where $j_{1},\dots,j_{k-1}$ are distinct indices, such that each of them corresponds to distinct $\gamma_{j_{k-1}}>\gamma_{j_{k-2}}>\dots>\gamma_{j_{1}}>\gamma_{1}$. We again represent $g$ as
$g=\tilde{g}\cdot(1,j_{1})\cdots(1,j_{k-2})(1,j_{k-1}),$
and arrive at a generalization of (18):
(25) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=N\mathrm{Trace}_{V_{\lambda}}\Bigl{(}\pi_{\tau}(T_{1})^{m}\Bigr{)}=\frac{\dim\tau}{(N-1)!}\sum_{k=1}^{m+1}(-\theta)^{k-1}\\ \times\sum_{\gamma}\sum_{j_{1},\dots,j_{k-1}}\sum_{\tilde{g}\in S_{N}\mid\tilde{g}\gamma=\gamma}\chi^{\tau}\bigl{(}\tilde{g}\cdot(1,j_{1})\cdots(1,j_{k-2})(1,j_{k-1})\bigr{)}p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]$
What remains is to carefully match the last formula with the claimed (23). First, the summation index $k$ runs from $1$ to $m+1$ in (25), however, for $k>p$, it is impossible to choose $k$ distinct $\gamma_{j_{k-1}}>\gamma_{j_{k-2}}>\dots>\gamma_{j_{1}}>\gamma_{1}$ and the sum is empty; hence, we can restrict the upper bound to be $\min(k+1,p)$.
Next, the summation over all $\gamma$ involves
$\frac{N!}{(n_{1})!(n_{2})!\cdots(n_{p})!}$
terms. Absorbing the factorials from the denominator into the character summation (they are present in (21)) and cancelling $N!$ with $(N-1)!$ in denominator in (25), we are left with a sum involving a single fixed $\gamma$ and $\pi_{\tau}\sum_{i=1}^{N}(T_{i})^{m}$.
We claim that the summation over $\tilde{g}$ in (25) gives the character averages of (21). Indeed, note that
$(1,j_{1})\cdots(1,j_{k-2})(1,j_{k-1})=(j_{k-1},j_{k-2},\dots,j_{1},1)$
and the corresponding degrees $\gamma_{j_{k-1}},\gamma_{j_{k-2}},\dots,\gamma_{1}$ are arranged in the decreasing order, which correspond to increasing indices of $d_{i}$ — matching the cycle $\mathfrak{c}$ used in (21). The stabilizer of $\gamma$, i.e. the subgroup $\tilde{g}\in S_{N}\mid\tilde{g}\gamma=\gamma$ is not necessary $S_{n_{1}}\times S_{n_{2}}\times\dots S_{n_{p}}$, but it is isomorphic to this group, with isomorphism obtained by renaming $\{1,2,\dots,N\}$. Since the characters are central, this leads to the same function as (21).
It remains to identify the sums of $p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]$. Note that when we act with $(T_{i})^{m}$ using its definition (9), the factors $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ correspond to various ~$\tilde{\ell}_{a}$. On the other hand, each application of a term from $\theta\sum_{j\mid\gamma_{j}>\gamma_{i}}(i,j)x^{\gamma}$, leads to addition of an additional index ${j_{a}}$. In this index, only the value of $\gamma_{j_{a}}$ matters. The set $\mathcal{A}$ in (23) is identified with $\{\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}\}$ in (25), leading to the combinatorial factor $\prod_{a\in\mathcal{A}}n_{a}$, counting the number of ways to choose indices $\gamma_{j_{a}}$ corresponding to $\mathcal{A}$ (the $n_{a}$ corresponding to $\gamma_{1}$ should be also accounted, as it arises from the first application of one of the operators $T_{i}$ in $\sum_{i=1}^{N}(T_{i})^{m}$). Once the set $\mathcal{A}$ is fixed, the sums of products of ~$\tilde{\ell}_{a}$ coming from $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ in $T_{i}$ combine precisely into $h^{\mathcal{A}}_{m+1-k}$ of (23). ∎
Lemma 5.5.
For the trivial character, $\tau=(N)$, the formulas of Theorems 5.4 and 3.1 are the same.
Proof.
The formula (5) can be rewritten in terms of the $m$–th power of $N\times N$ matrix as
(26) $\begin{pmatrix}1&1&1&\ldots&1\end{pmatrix}\begin{pmatrix}\ell_{1}&-\theta&-\theta&\ldots&-\theta\\ 0&\ell_{2}&-\theta&\ldots&-\theta\\ 0&0&\ell_{3}&\ldots&-\theta\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&\ell_{N}\end{pmatrix}^{m}\begin{pmatrix}1\\ 1\\ 1\\ \vdots\\ 1\end{pmatrix},$
where $\ell_{i}=\lambda_{i}+\theta(N-i)$. On the other hand, the formula (23) can rewritten in terms of $m$–th power of $p\times p$ matrix as
(27) $\begin{pmatrix}n_{1}&n_{2}&n_{3}&\ldots&n_{p}\end{pmatrix}\begin{pmatrix}\tilde{\ell}_{1}&-n_{2}\theta&-n_{3}\theta&\ldots&-n_{p}\theta\\ 0&\tilde{\ell}_{2}&-n_{3}\theta&\ldots&-n_{p}\theta\\ 0&0&\tilde{\ell}_{3}&\ldots&-n_{p}\theta\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&\tilde{\ell}_{p}\end{pmatrix}^{m}\begin{pmatrix}1\\ 1\\ 1\\ \vdots\\ 1\end{pmatrix},$
where ~$\tilde{\ell}_{i}$ are a subset of $\ell_{i}$ given by ~$\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p})$ in the notation (20). Our task is to prove (26)=(27). One can do this directly, but our approach is to proceed through generating functions. On the (26) side, the generating function was computed in Proposition 3.2. Let $\widetilde{\mathrm{eig}}_{m}(\lambda)$ denote the eigenvalues computed by (27). Repeating the argument of Proposition 3.2, we have
$1-\theta z\sum_{m=0}^{\infty}\widetilde{\mathrm{eig}}_{m}(\lambda)z^{m}=\sum_{\begin{smallmatrix}\mathcal{A}\subset\{1,\dots,p\}\end{smallmatrix}}(-\theta z)^{|\mathcal{A}|}\prod_{a\in\mathcal{A}}\frac{n_{a}}{1-\tilde{\ell}_{a}z}=\prod_{i=1}^{p}\left(1-\frac{n_{i}\theta z}{1-\tilde{\ell}_{i}z}\right)\\ =\prod_{i=1}^{p}\frac{1-(\tilde{\ell}_{i}+n_{i}\theta)z}{1-\tilde{\ell}_{i}z}.$
The last expression matches the right-hand side of (7), once we notice the telescoping cancellations in (7). ∎
5.3. $(N-1,1)$ isotype
While the formula of Theorem 5.4 is very general, it produces an answer in terms of the functions $\chi^{\tau}[\mathcal{A};\mathbf{n}]$. Evaluation of these functions is a separate task. As we have already mentioned right after Theorem 5.4, for $\tau=(N)$ and $\tau=1^{N}$ the evaluation is particularly simple. Another case where it is fully explicit is $\tau=(N-1,1)$.
Theorem 5.6.
Let $\tau=(N-1,1)$ and $\mathbf{n}=(n_{1},n_{2},\dots,n_{p}>0)$. Then we have
(28) $\chi^{\tau}[\mathcal{A};\mathbf{n}]=p-1-\sum_{a\in\mathcal{A}}\frac{1}{n_{a}}.$
Proof.
The irreducible representation of $S_{N}$ of type $\tau=(N-1,1)$ is the standard $(N-1)$–dimensional representation in the space of vectors $(x_{1},\dots,x_{N})$ with $\sum_{i=1}^{N}x_{i}=0$ and action $\sigma\circ(x_{1},\dots,x_{N})=(x_{\sigma^{-1}(1)},\dots,x_{\sigma^{-1}(N)})$. Evaluating the character as trace, we get an expression in terms of the number of the fixed points:
$\chi^{\tau}(\sigma)=\#\{1\leq i\leq N\mid\sigma(i)=i\}-1,\qquad\tau=(N-1,1).$
Recalling the definition (21), we need to compute
(29) $\sum_{\tilde{g}\in S_{n_{1}}\times S_{n_{2}}\times\dots\times S_{n_{p}}}\,\,\chi^{\tau}(\tilde{g}\cdot\mathfrak{c})=\sum_{\tilde{g}}\sum_{i}\mathbf{1}_{\tilde{g}\cdot\mathfrak{c}(i)=i}-\prod_{j=1}^{p}(n_{j})!,$
where $\mathfrak{c}$ is a $|\mathcal{A}|$-cycle permuting the groups correspinding to the elements $a\in\mathcal{A}$. Changing the order of summation in (29), we get
(30) $\sum_{q=1}^{p}\,\sum_{i=n_{1}+\dots+n_{q-1}+1}^{n_{1}+\dots+n_{q}}\,\sum_{\tilde{g}\in S_{n_{1}}\times S_{n_{2}}\times\dots\times S_{n_{p}}}\mathbf{1}_{\tilde{g}\cdot\mathfrak{c}(i)=i}-\prod_{j=1}^{p}(n_{j})!\\ =\sum_{q=1}^{p}\left[\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)+\mathbf{1}_{q\not\in\mathcal{A}}n_{q}\right](n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right]-\prod_{j=1}^{p}(n_{j})!,\\ =(p-1)\prod_{j=1}^{p}(n_{j})!-\sum_{q=1}^{p}\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right].$
Dividing by $\prod\limits_{j=1}^{p}(n_{j})!$, we get (28). ∎
5.4. $a\dots ab\dots b$ case with $p=2$
Another situation where the general formula of Theorem 5.4 is fully explicit is $p=2$. In fact, in this special case we are able to compute all eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ rather than only their sums over $i$. We fix $N=1,2,\dots$ and $0\leq\eta\leq N$ and set the multiplicities to be $(n_{1},n_{2})=(N-\eta,\eta)$. We further choose $a>b\geq 0$ and set
(31) $\lambda=\underbrace{a,a,\dots,a}_{N-\eta},\underbrace{b,b,\dots,b}_{\eta}.$
Corollary 5.7.
For $\lambda$ in (31) and $\tau=(N-k,k)$ with $0\leq k\leq\min(\eta,N-\eta)$, we have
$\mathrm{eig}_{m}(\lambda,\tau,i)=(a+\theta\eta)^{m}(N-\eta)+b^{m}\eta-\theta\frac{(a+\theta\eta)^{m}-b^{m}}{a+\theta\eta-b}\bigl{(}\eta(N-\eta)-k(N-k+1)\bigr{)},$
for all $1\leq i\leq\dim V_{\lambda,\tau}$. For other $\tau$, the space $V_{\lambda,\tau}$ is empty.
Proof.
We need some facts from the representation theory of symmetric groups, which can be found e.g. in [Dun78], or [Sta84, Section 3], or [CSST08, Sections 6.1, 6.2]. In particular, $V_{\lambda}$ expands into a direct multiplicity one sum of irreducible representations $\tau=(N-k,k)$, $0\leq k\leq\min(\eta,N-\eta)$. The operator $(T_{1}^{m}+\dots+T_{N}^{M})$ commutes with the action of symmetric group. Therefore, by the Schur’s lemma, it necessary acts as a multiple of identity in each $V_{\lambda;\tau}$ (which would not have been true if $\tau$ had multiplicity larger than $1$ in $\lambda$). Hence, the eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ do not depend on $i$ and we can get them by dividing (23) by $\dim\tau$.
Let us evaluate the ingredients of (23). $(S_{N},S_{N-\eta}\times S_{\eta})$ is a Gelfand pair, which means that each irreducible representation of $S_{N}$ has at most one $S_{N-\eta}\times S_{\eta}$–invariant vector. Interpreting summations over $\tilde{g}\in S_{N-\eta}\times S_{\eta}$ in (21) as projections on invariants, we identify $\chi^{\tau}[\mathcal{A};\mathbf{n}]$ with spherical functions corresponding to this invariant vector. The latter spherical function is evaluated in [Dun78, Corollary 2.2], [Sta84, (3.15)], [CSST08, Theorems 6.1.10, 6.2.3] in terms of hypergeometric function ${}_{3}F_{2}$, which can also be identified with the Hahn polynomial. The value at length $2$ cycle is given as
(32) ${}_{3}F_{2}\left(\begin{smallmatrix}-k,\,k-N-1,\,-1\\ \eta-N,-\eta\end{smallmatrix}\Bigr{|}1\right)=1+\frac{(-k)(k-N-1)(-1)}{(-\eta)(\eta-N)1!}=1-\frac{k(N-k+1)}{\eta(N-\eta)}.$
Hence, (23) specializes to:
$\sum_{\begin{smallmatrix}\mathcal{A}\subset\{1,2\}\\ |\mathcal{A}|=1\end{smallmatrix}}h^{\mathcal{A}}_{m}\cdot\prod_{\alpha\in\mathcal{A}}n_{\alpha}-\theta\sum_{\begin{smallmatrix}\mathcal{A}\subset\{1,2\}\\ |\mathcal{A}|=2\end{smallmatrix}}h^{\mathcal{A}}_{m-1}\cdot\left(1-\frac{k}{\eta}\cdot\frac{N-\eta+1}{N-\eta}\right)\cdot\prod_{\alpha\in\mathcal{A}}n_{\alpha}.\\ =(a+\theta\eta)^{m}(N-\eta)+b^{m}\eta-\theta\frac{(a+\theta\eta)^{m}-b^{m}}{a+\theta\eta-b}\left(1-\frac{k}{\eta}\cdot\frac{N-k+1}{N-\eta}\right)\eta(N-\eta).\qed$
6. Example: Eigenvalues for $N=3$.
Here is the full list of the eigenvalues of the operator $\mathcal{P}_{m}$ for $N=3$. We keep using the notation (3). When we speak about leading monomials in the description of eigenfunctions, we use the point of view of Theorem 4.3 and describe the part of $F_{\lambda,\tau,i}$ in $V_{\lambda}$.
  1. (1)
    $\lambda=(a,a,a)$: For each $a\geq 0,$ $\mathcal{P}_{m}$ has an eigenvalue $3a^{2}m$ on the eigenfunction $x_{1}^{a}x_{2}^{a}x_{3}^{a}$.
  2. (2)
    $\lambda=(a,a,b)$: For each $a>b\geq 0$, $\mathcal{P}_{m}$ has:
    1. (a)
      Eigenvalue $2(a+\theta)^{m}+b^{m}-2\theta\frac{(a+\theta)^{m}-b^{m}}{a+\theta-b}$ on the eigenfunction with leading monomials $x_{1}^{a}x_{2}^{a}x_{3}^{b}+x_{1}^{a}x_{2}^{b}x_{3}^{a}+x_{1}^{b}x_{2}^{a}x_{3}^{a}$.
    2. (b)
      Two equal eigenvalues $2(a+\theta)^{m}+b^{m}+\theta\frac{(a+\theta)^{m}-b^{m}}{a+\theta-b}$ on two eigenfunctions with leading monomials in the space spanned by $x_{1}^{a}x_{2}^{a}x_{3}^{b}-x_{1}^{a}x_{2}^{b}x_{3}^{a}$ and $x_{1}^{a}x_{2}^{a}x_{3}^{b}-x_{1}^{b}x_{2}^{a}x_{3}^{a}$.
  3. (3)
    $\lambda=(a,b,b)$: For each $a>b\geq 0$, $\mathcal{P}_{m}$ has:
    1. (a)
      Eigenvalue $(a+2\theta)^{m}+2b^{m}-2\theta\frac{(a+2\theta)^{m}-b^{m}}{a+2\theta-b}$ on the eigenfunction with leading monomials $x_{1}^{a}x_{2}^{b}x_{3}^{b}+x_{1}^{b}x_{2}^{a}x_{3}^{b}+x_{1}^{b}x_{2}^{b}x_{3}^{a}$.
    2. (b)
      Two equal eigenvalues $(a+2\theta)^{m}+2b^{m}+\theta\frac{(a+2\theta)^{m}-b^{m}}{a+2\theta-b}$ on two eigenfunctions with leading monomials in the space spanned by $x_{1}^{a}x_{2}^{b}x_{3}^{b}-x_{1}^{b}x_{2}^{a}x_{3}^{b}$ and $x_{1}^{a}x_{2}^{b}x_{3}^{b}-x_{1}^{b}x_{2}^{b}x_{3}^{a}$.
  4. (4)
    $\lambda=(a,b,c)$: For each $a>b>c\geq 0$, $\mathcal{P}_{m}$ has (we denote $(\ell_{1},\ell_{2},\ell_{3})=(a+2\theta,b+\theta,c)$):
    1. (a)
      Eigenvalue $h^{(1)}_{m}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}-\theta h^{(2)}_{m-1}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}+\theta^{2}h^{(3)}_{m-2}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}$ on the eigenfunction with leading monomials
      $x_{1}^{a}x_{2}^{b}x_{3}^{c}+x_{1}^{b}x_{2}^{a}x_{3}^{c}+x_{1}^{a}x_{2}^{c}x_{3}^{b}+x_{1}^{c}x_{2}^{b}x_{3}^{a}+x_{1}^{b}x_{2}^{c}x_{3}^{a}+x_{1}^{c}x_{2}^{a}x_{3}^{b}.$
    2. (b)
      Two pairs of equal eigenvalues
      (33) $\left(\ell_{1}^{m}+\ell_{2}^{m}+\ell_{3}^{m}\right)-\frac{\theta^{2}}{2}h_{m-2}\left(\ell_{1},\ell_{2},\ell_{3}\right)\\ \pm\frac{\theta}{2}h_{m-2}\left(\ell_{1},\ell_{2},\ell_{3}\right)\sqrt{4\left(\ell_{1}^{2}+\ell_{2}^{2}+\ell_{3}^{2}\right)-4\left(\ell_{1}\ell_{2}+\ell_{1}\ell_{3}+\ell_{2}\ell_{3}\right)-3\theta^{2}}$
      on four eigenfunctions with leading monomials in the space spanned by
      (34) $x_{1}^{a}x_{2}^{b}x_{3}^{c}-x_{1}^{b}x_{2}^{c}x_{3}^{a},\quad x_{1}^{a}x_{2}^{b}x_{3}^{c}-x_{1}^{c}x_{2}^{a}x_{3}^{b},\quad x_{1}^{b}x_{2}^{a}x_{3}^{c}-x_{1}^{a}x_{2}^{c}x_{3}^{b},\quad x_{1}^{b}x_{2}^{a}x_{3}^{c}-x_{1}^{c}x_{2}^{b}x_{3}^{a}.$
    3. (c)
      Eigenvalue $h^{(1)}_{m}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}+\theta h^{(2)}_{m-1}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}+\theta^{2}h^{(3)}_{m-2}\bigl{(}\ell_{1},\ell_{2},\ell_{3}\bigr{)}$ on the eigenfunction with leading monomials
      $x_{1}^{a}x_{2}^{b}x_{3}^{c}-x_{1}^{b}x_{2}^{a}x_{3}^{c}-x_{1}^{a}x_{2}^{c}x_{3}^{b}-x_{1}^{c}x_{2}^{b}x_{3}^{a}+x_{1}^{b}x_{2}^{c}x_{3}^{a}+x_{1}^{c}x_{2}^{a}x_{3}^{b}.$
(1a), (2a), (3a), and (4a) are all instances of Theorem 3.1 or Theorem 5.4 with $\tau=(3)$. (2ab) and (3ab) are instances of Corollary 5.7. (4a) and (4c) are instances of Theorem 5.1 for $\tau=(3)$ and $\tau=(1,1,1)$, respectively. The sum of all four eigenvalues in (4b) matches Theorem 5.1 for $\tau=(2,1)$, where we have $\dim\tau=\chi^{\tau}(\mathrm{Id})=2$, $\chi^{\tau}\bigl{(}(1,2)\bigr{)}=0$, $\chi^{\tau}\bigl{(}(1,2,3)\bigr{)}=-1$. The individual eigenvalues in (4b) are not computed by any of our theorems; the formula (33) can be obtained by evaluating the matrices of the operators $T_{1}^{m}$, $T_{2}^{m}$, $T_{3}^{m}$ by diagonalization and then direct computation of the eigenvalues of $T_{1}^{m}+T_{2}^{m}+T_{3}^{m}$ (we can either deal with $6\times 6$ matrices in the space of all permutations of $x_{1}^{a}x_{2}^{b}x_{3}^{c}$ or $4\times 4$ matrices in the subspace spanned by (34)).
7. Appendix: Basic properties of $\mathcal{P}_{m}$
In this section we discuss the basic properties of the operators $\mathcal{P}_{m}$, mentioned in Section 2. These properties are well-known, we provide proofs only for the sake of being self-contained.
Lemma 7.1.
Operators $\mathcal{P}_{m}$, $m=1,2,\dots$, commute.
Proof.
We develop a table of commutators. We start by splitting the Dunkl operator of Definition 2.1 into two parts and write $\mathcal{D}_{i}=\frac{\partial}{\partial x_{i}}+\theta\Delta_{i}$ with $\Delta_{i}=\sum_{j\neq i}\frac{1-(i,j)}{x_{i}-x_{j}}$. We claim
(35) $\Delta_{a}\Delta_{b}=\Delta_{b}\Delta_{a},\qquad 1\leq a,b\leq N.$
In order to prove (35) for $a\neq b$, plug the definition of $\Delta_{a}$ into $\Delta_{a}\Delta_{b}$, open the parenthesis and note that the double sum splits into three types of summands: involving two, three, and four variables $x_{i}$. The parts involving two and four variables are obviously the same in $\Delta_{a}\Delta_{b}$ and in $\Delta_{b}\Delta_{a}$. The part involving three variables, $x_{a}$, $x_{b}$, $x_{j}$ is more tricky, it is:
(36) $\frac{1-(a,j)}{x_{a}-x_{j}}\cdot\frac{1-(b,j)}{x_{b}-x_{j}}+\frac{1-(a,b)}{x_{a}-x_{b}}\cdot\frac{1-(b,j)}{x_{b}-x_{j}}+\frac{1-(a,j)}{x_{a}-x_{j}}\cdot\frac{1-(b,a)}{x_{b}-x_{a}},$
where in each fraction we first apply the operator in numerator and then divide by the denominator. The desired relation (35) follows from the symmetry of (36) in $a\leftrightarrow b$, which is verified directly by moving all permutation operators to the right. The next useful identity, which is verified directly, is that for $a\neq b$
(37) $[x_{a},\Delta_{b}]=x_{a}\Delta_{b}-\Delta_{b}x_{a}=(a,b).$
Another direct verification yields that for $a\neq b$ we have
(38) $\left[\frac{\partial}{\partial x_{a}},\Delta_{b}\right]=\frac{1-(a,b)}{(x_{a}-x_{b})^{2}}+\frac{1}{x_{a}-x_{b}}\left(\frac{\partial}{\partial x_{a}}-\frac{\partial}{\partial x_{b}}\right)(a,b).$
The last identity makes it clear that $\left[\tfrac{\partial}{\partial x_{a}},\Delta_{b}\right]=\left[\tfrac{\partial}{\partial x_{b}},\Delta_{a}\right]$. Combining with (35), this implies commutativity of the Dunkl operators: $\mathcal{D}_{a}\mathcal{D}_{b}=\mathcal{D}_{b}\mathcal{D}_{a}$. Next, for each $a\neq b$ we have
(39) $[x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}]=\theta(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b).$
Indeed, using (37) and commutativity of $x_{a}$ with $\frac{\partial}{\partial x_{b}}$, we have
$x_{a}\mathcal{D}_{a}x_{b}\mathcal{D}_{b}-x_{b}\mathcal{D}_{b}x_{a}\mathcal{D}_{a}=x_{a}x_{b}\mathcal{D}_{a}\mathcal{D}_{b}-x_{a}(a,b)\mathcal{D}_{b}-x_{b}x_{a}\mathcal{D}_{b}\mathcal{D}_{a}+\theta x_{b}(a,b)\mathcal{D}_{a}.$
Cancelling the first and third terms and using $(a,b)\cdot\mathcal{D}_{b}\cdot(a,b)=\mathcal{D}_{a}$ we arrive at (39). Using this identity we further compute
(40) $[x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}[x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}](x_{b}\mathcal{D}_{b})^{m-k-1}\\ =\theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b)(x_{b}\mathcal{D}_{b})^{m-k-1}=\\ \theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(x_{a}\mathcal{D}_{a})^{m-k-1}\cdot(a,b)=\theta\bigl{(}(x_{a}\mathcal{D}_{a})^{m}-(x_{b}\mathcal{D}_{b})^{m}\bigr{)}\cdot(a,b).$
Iterating the same argument again, we get
$[(x_{a}\mathcal{D}_{a})^{l},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}[x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}](x_{a}\mathcal{D}_{a})^{l-k-1}\\ =\theta\left(\sum_{k=m}^{m+l-1}(x_{a}\mathcal{D}_{a})^{m+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
Interchanging $l$ and $m$ and using the skew-symmetry $[f,g]=-[g,f]$, we also have
$[(x_{b}\mathcal{D}_{b})^{l},(x_{a}\mathcal{D}_{a})^{m}]=\theta\left(\sum_{k=0}^{m-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=l}^{m+l-1}(x_{a}\mathcal{D}_{a})^{l+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
Summing the last two identities over all $a<b$, we arrive at
(41) $\left[\sum_{a=1}^{N}(x_{a}\mathcal{D}_{a})^{l},\sum_{b=1}^{N}(x_{b}\mathcal{D}_{b})^{m}\right]\\ =\theta\sum_{a<b}\left(\sum_{k=0}^{m-1}+\sum_{k=m}^{l+m-1}-\sum_{k=0}^{l-1}-\sum_{k=l}^{m+l-1}\right)(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-1-k}(a,b)=0.\qed$
Lemma 7.2.
The restriction of the operator $\mathcal{P}_{2}$ on the space of symmetric polynomials coincides with
(42) $\sum_{i=1}^{N}\left(x_{i}\frac{\partial}{\partial x_{i}}\right)^{2}+\theta\sum_{i\neq j}\frac{x_{i}(x_{i}+x_{j})}{x_{i}-x_{j}}\frac{\partial}{\partial x_{i}}.$
Remark 7.3.
Changing the variables $x_{a}=\exp(\mathbf{i}z_{a})$, conjugating with $\prod_{a<b}\sin^{\theta}\left(\frac{z_{a}-z_{b}}{2}\right)$ and shifting by a constant, (42) turns into $-H$, where $H$ is the trigonometric version of the so-called Calogero-Moser-Sutherland Hamiltonian:
$H=\sum_{i=1}^{N}\left(\frac{\partial}{\partial z_{i}}\right)^{2}+\frac{\theta(\theta-1)}{2}\sum_{i<j}\frac{1}{\sin^{2}\left(\frac{z_{a}-z_{b}}{2}\right)}.$
Proof of Lemma 7.2.
Since $01-(i,j)=$ 0 on symmetric functions, $(x_{i}\mathcal{D}_{i})^{2}$ restricts as
$\left(x_{i}\frac{\partial}{\partial x_{i}}+\theta x_{i}\sum_{j\neq i}\frac{1-(i,j)}{x_{i}-x_{j}}\right)\left(x_{i}\frac{\partial}{\partial x_{i}}\right)=\left(x_{i}\frac{\partial}{\partial x_{i}}\right)^{2}+\theta\sum_{j\neq i}\frac{x_{i}}{x_{i}-x_{j}}\left[x_{i}\frac{\partial}{\partial x_{i}}-x_{j}\frac{\partial}{\partial x_{j}}(i,j)\right].$
The permutation $(i,j)$ acts identically on symmetric functions and can be removed from the right-hand side of the last formula. Summing over all $i$, we arrive at (42). ∎
Lemma 7.4.
For identical representation $\tau=(N)$ the joint eigenfunction $F_{\lambda,(N),1}$ of operators $\mathcal{P}_{m}$ from Theorem 4.3 is (up to a constant factor) the symmetric Jack polynomial $J_{\lambda}(x_{1},\dots,x_{N};\,\theta)$.
Proof.
As in the proof of Theorem 3.1, we note that the linear space $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ is invariant both for $\mathcal{P}_{m}$ and for the symmetric group $S_{N}$. Because $S_{N}$–action commutes with $\mathcal{P}_{m}$ action, $\mathcal{P}_{m}$ preserves the subspace of symmetric polynomial inside $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. Hence, $\mathcal{P}_{m}$ must have some symmetric eigenfunctions in $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ and then by triangularity consideration, the unique symmetric eigenfunction with leading monomial in $V_{\lambda}$ has to be $F_{\lambda,(N),1}$. We conclude that $F_{\lambda,(N),1}$ is symmetric. Since $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, its eigenfunctions are orthogonal with respect to scalar product of that lemma. We conclude that as $\lambda$’s vary, the functions $F_{\lambda,(N),1}$ form an orthogonal basis of the space $\Lambda_{N}$ of symmetric functions in $N$ variables and this basis is related to monomial symmetric functions $m_{\lambda}$ by a triangular transformation. Being such an orthogonal basis is a property which uniquely determines Jack polynomials and can be taken as one of their equivalent definitions, see [Sta89] or [Mac95, Section 6.10] and note that they use $\alpha=1/\theta$ as the parameter. ∎